OpenStudy (anonymous):

Find a quadratic equation with roots -1+4i and -1-4i.

OpenStudy (samanthagreer):

x2−10x+29

OpenStudy (anonymous):

Well, if those were the roots, it would factor as: \[ (x-(-1+4i))(x-(-1-4i)) = 0 \]

OpenStudy (anonymous):

So I say foil.

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

\[ x^2 - (-1+4i)x - (-1-4i)x + (-1+4i) (-1-4i) \]

OpenStudy (anonymous):

is x^2-10x+29 the answer?

OpenStudy (anonymous):

one way to do it is to remember that the quadratic with zeros of \(a\pm bi\) is \[x^2-2ax+a^2+b^2\]

OpenStudy (anonymous):

so no, your answer is incorrect

OpenStudy (anonymous):

another way is to set \[x=-1+4i\] and work backwards add \(1\) \[x+1=4i\]then square (carefully)\[(x+1)^2=(i)^2\\ x^2+2x+1=-16\]

OpenStudy (anonymous):

thank you so much!